OpenStudy (anonymous):

Why would f(1/4) be undefined?

OpenStudy (anikate):

well we no that the answer is f/4

OpenStudy (anikate):

there is no equal sign

OpenStudy (asnaseer):

your question is not complete. you need to tel us what f(x) is defined to be.

OpenStudy (anonymous):

I have to explain why why f(1) = 3, f(0) = 2, and f(-1) = 1, but why f(1/4) is defined

OpenStudy (anonymous):

*undefined

OpenStudy (asnaseer):

have you been given the expression for f(x) in terms of x?

OpenStudy (anonymous):

in the 4x^2+ 7x - 2/ 4x-1

OpenStudy (asnaseer):

ok - now it makes sense

OpenStudy (anonymous):

i simplified it to x + 2/ 1 XD

OpenStudy (asnaseer):

if you look at the last term, you have a division by "4x-1". what will "4x-1" equate to when x=1/4?

OpenStudy (asnaseer):

\[f(x)=4x^2+7x-\frac{2}{4x-1}\]

OpenStudy (anonymous):

ohhh it would be 0

OpenStudy (asnaseer):

correct - so you would be dividing by zero - which is undefined.

OpenStudy (asnaseer):

BTW: your simplification to x + 2/ 1 is not correct.

OpenStudy (anonymous):

ahhhh :( what I do wrong?

OpenStudy (asnaseer):

can you please try and explain to me step by step what you did so that I can spot where you went wrong

OpenStudy (anonymous):

I factored to get (x+2)(4x-1)/4x-1 then canceled out common factors and got x+2/1

OpenStudy (asnaseer):

your factorisation in wrong. remember the last term is a fraction.

OpenStudy (asnaseer):

\[\begin{align} f(x)&=4x^2+7x-\frac{2}{4x-1}\\ &=\frac{(4x-1)(4x^2+7x)-2}{4x-1} \end{align}\]

OpenStudy (anonymous):

its supposed to be a rational expression o.O

OpenStudy (anonymous):

it looks like this : 4x^2+ 7x - 2 -------------- 4x-1

OpenStudy (asnaseer):

so the equation is supposed to be:\[f(x)=\frac{4x^2+7x-2}{4x-1}\]

OpenStudy (asnaseer):

which can be factored as:\[f(x)=\frac{(x+2)(4x-1)}{4x-1}\]

OpenStudy (anonymous):

yes :)

OpenStudy (asnaseer):

but you cannot cancel the 4x-1 because it would be like dividing top and bottom by zero when x=1/4

OpenStudy (anonymous):

i just canceled out common factors that what my teacher told me :(

OpenStudy (asnaseer):

you can only do that if you are told that x can NEVER be equal to 1/4

OpenStudy (asnaseer):

in this case I believe x can take on ay real value from minus infinity to plus infinity

OpenStudy (asnaseer):

and it is undefined at x=1/4

OpenStudy (asnaseer):

was the question asking you why is this undefined at x=1/4?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

so then you can state that it undefined at x=1/4 because at x=1/4 we would be dividing by zero.

OpenStudy (anonymous):

okay Thank you!!! XD Wish i could giv you more than one medal o.O lol

OpenStudy (asnaseer):

You're welcome - I'm glad I was able to explain it well enough for you.